OpenStudy (anonymous):

An apple orchard is selling apple cider in spherical sipping containers. If each container is 7 cm in radius, what volume of liquid is in each container? Round to the nearest cubic centimeter.

OpenStudy (turingtest):

what is the formula for the volume of a sphere?

OpenStudy (anonymous):

V=4/3πr3 I use it but i got 808 and its wong

OpenStudy (anonymous):

v=4/3 pie r^3

OpenStudy (turingtest):

i think you are plugging the numbers into your calculator wrong first do 7^3, then multiply by 4, then divide by 3, then multiply by pi

OpenStudy (anonymous):

@maryroll3 First find the value of r^3. Since we know the radius is 7, what is 7^3?

OpenStudy (anonymous):

Btw, this is AnswerMyQuestions here. :P

OpenStudy (anonymous):

@maryroll3 Do you know 7^3?

OpenStudy (anonymous):

nvm, i solve it already Volume = (4/3) (3.142)(7) 3 Volume = 1436.76